x+y+3z+2w=13
x-2y+2z+w=8
3x+y+z-2w=1

Answer by Alan3354(69443)

x+y+3z+2w=13
x-2y+2z+w=8
3x+y+z-2w=1
-------------
4 variables, 3 equations.
What do you want to do?